Problem 5-5 Calculating the Number of Periods [LO4] Solve for the unknown number of years in each of the following (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.):Present Value Years Interest Rate Future Value $490 10 % $1,252 740 11 1,670 17,700 16 216,050 20,800 13 262,330
Answer #1

t = ln(FV / PV) / ln(1 + r)

t = ln($1,252 / $490) / ln 1.10 = 9.84 years

t = ln($1,670 / $740) / ln 1.11 = 7.80 years

t = ln($216,050 / $17,700) / ln 1.16 = 16.86 years

t = ln($262,330 / $20,800) / ln 1.13 = 20.74 years